OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (5,4): m = -3

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

point-slope form is y - y1 = m(x - x1) where y1 is the y part of the point, x1 is the x part of the point, and m is the slope. so just plug all that in

OpenStudy (anonymous):

Y = -3 (x - 5) ?

OpenStudy (aripotta):

y - 4 = -3(x - 5)

OpenStudy (anonymous):

How did you get y - 4 ?

OpenStudy (aripotta):

4 is the y part of the point (5,4) and the formula is y - y1 = m(x - x1). so y - 4

OpenStudy (anonymous):

Ohh, okay. Thank you! :)

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\)